OpenStudy (drewbooks):

If 2i + (x + yi) = 6, what is x + yi?

myininaya (myininaya):

collect your imaginary parts and your real parts then compare both sides

OpenStudy (drewbooks):

What do I do with the y?

myininaya (myininaya):

for example if ai+(b+ci)=e+di b+ai+ci=e+di b+(a+c)i=e+di means b=e and a+c=d

myininaya (myininaya):

you have 2i+(x+yi) on your left hand side 6+0i on the right hand side

myininaya (myininaya):

can you identify the real part on both sides of your equation? can you identify the imaginary part on both sides of your equation?

OpenStudy (drewbooks):

2i, yi, and 0i are imaginary, x and 6 are real?

myininaya (myininaya):

so x is 6 then

myininaya (myininaya):

and 2i+yi=0i

myininaya (myininaya):

factor out the i on the left hand expression of 2i+yi=0i

OpenStudy (drewbooks):

y is -2

OpenStudy (drewbooks):

So 6 -2i is my answer?

myininaya (myininaya):

yes (2+y)i=0i implies 2+y =0 so y=-2

myininaya (myininaya):

yes

OpenStudy (drewbooks):

Awesome, thank you.

myininaya (myininaya):

np